OpenStudy (anonymous):

find the quotient. 45 divided by 5/11 (p.s i suck at math)

OpenStudy (hba):

\[45 \div \frac{ 9 }{ 5 }\] Right ?

OpenStudy (hba):

Use this rule. \[a \div \frac{ b }{ c }=a \times \frac{ c }{ b }\]

OpenStudy (hba):

^ good to hear :)

OpenStudy (hba):

@ErinSlack01 I guess it's a he though. Start working.

OpenStudy (texaschic101):

45 / (5/11) When dividing fractions....you " flip " the number you are dividing by, and multiply. 45/1 x 11/5 = your answer

OpenStudy (hba):

@Hero Not a girly name you see :/

OpenStudy (anonymous):

im a gir

OpenStudy (texaschic101):

Erin can be a girls name...but it can also be a boys name...

OpenStudy (hba):

@Hero Well,I don't know. I just don't know much because i live in another country where these kind of names aren't used :P

OpenStudy (texaschic101):

what about the other 1%....lol

OpenStudy (hba):

Well i check the dictionary, A female given name. Erin is a female who has the body of a goddess and the brain of a genius

OpenStudy (hba):

@ErinSlack01 Do you agree with this ?

OpenStudy (texaschic101):

Its a Gaelic origin, and has been associated with Ireland.